Example 1 An antibiotic suspension contains 250 milligrams of drug in 1 teaspoon of medication. How many grams of drug are coPharmacy Calculations please use conversion factors as well as show the bridges between metric and imperial system if possible. Thanks so much!
Answer #1

1) concentration of the suspension = 250 milligrams drug in 1 teaspoon medication.

Two teaspoons contain double the medication. Hence it contains double the drug.

which is 2*250 = 500 milligrams.

5) Weight of potassium chloride dissolved = 4 grams

Final volume of solution = 25 millilitres

As solute and solution are given in grams and millilitres, concentration percentage will be in (w/v)%

Percentage concentration of solution is (weight of KCl in grams/volume of solution in millilitres)*100

which is (4/25)*100 = 16% w/v

4) Volume and concentration of given solution is 100ml and 5%.

Let us suppose it is (w/v)%. Then 5 grams of solute is present in 100 ml solution.

Required concentration is 2% and available solute is 5 grams

Then the volume of solution is given by (solute in grams/percentage)*100

which is (5/2)*100 = 250 ml

250 millilitres of 2% solution can be prepared.

6) Volume or glycerin (solute) is 90 millilitres. Volume of water (solvent) is 600 millilitres.

Volume or solution is ( volume of solute + volume or solvent ) = 90 + 600 which is 690ml

The percentage v/v of solution is (volume or solute in ml / volume of solvent) * 100

which is (90/690)*100 = 13%v/v

7) W/W concentration of ointment is 0.002 % beclomethasone. This means 0.002 grams of beclomethasone is present in 100 grams of drug.

Given weight of ointment is 2 ounces which is 28.35*2 = 56.7 grams

Weight of beclomethasone in this ointment ( percentage * weight of ointment ) / 100 grams

which is ( 0.002 * 56.7 ) / 100 grams = 0.1134 grams or 113.4 milligrams

The weight of beclomethasone in 2 ounces of ointment is 113.4 milligrams

This is a brief explanation for the given calculations.. Hope this helps you out. Please feel free to give a like and contact us if you need further detail explanation. Thank you and have a great day.